# industrial control system STIX Structured Threat Information Expression

### v1

@@ 5 Analysis of Vulnerability Information Sharing
In addition to studying cyber threat activity, the analysis also reviews the sharing of vulnerability advisories associated with known exploitable ICS vulnerabilities based on their inclusion in the CISA KEV Catalog. This list was produced by reviewing the KEV Catalog for Common Vulnerabilities and Exposuress identified since 2023, which are associated with CISA ICS Alerts. Since these vulnerabilities have been known to be exploited by threat actors, they present a clear risk to asset owner environments, requiring some mitigating actions. While within IT environments these vulnerabilities are typically patched, it is well documented asset owners struggle to quickly deploy patches [sans_ics_ot_cybersecurity]. If an asset owner cannot quickly patch these vulnerabilities to mitigate their risk, they may need to deploy detection mechanisms to prevent their exploitation. Therefore, vulnerability information is reviewed in accordance with the same requirements as defined in Section 2. That is, evaluating whether the vendor provided some rule to support the detection of the vulnerability, or whether sufficient technical details are available about the product and associated vulnerability for the asset owner to generate their own detection.
@@ 6 Challenges in Threat Information Sharing
    • Integrity checks: Further, many controllers also use integrity check mechanisms, typically based on CRCs or Checksums, to verify that programs have not changed  [plc_security_practices]. On the one hand, security products can verify that a specific program has been downloaded to the device by validating its signatures. On the other hand, CRC/Checksum integrity checks are not sufficient against threat actors who intentionally modify the PLC code  [schuett2014evaluation]. There is a need for a wider deployment of secured integrity signatures  [banerjee2024designing].

### v2

@@ 5 Analysis of Vulnerability Information Sharing
In addition to studying cyber threat activity, the analysis also reviews the sharing of vulnerability advisories associated with known exploitable ICS vulnerabilities based on their inclusion in the CISA KEV Catalog. This list was produced by reviewing the KEV Catalog for Common Vulnerabilities and Exposuress identified since 2023, which are associated with CISA ICS Alerts. Since these vulnerabilities have been known to be exploited by threat actors, they present a clear risk to asset owner environments, requiring some mitigating actions. While within IT environments these vulnerabilities are typically patched, it is well documented asset owners struggle to quickly deploy patches [28]. If an asset owner cannot quickly patch these vulnerabilities to mitigate their risk, they may need to deploy detection mechanisms to prevent their exploitation. Therefore, vulnerability information is reviewed in accordance with the same requirements as defined in Section 2. That is, evaluating whether the vendor provided some rule to support the detection of the vulnerability, or whether sufficient technical details are available about the product and associated vulnerability for the asset owner to generate their own detection.
@@ 6 Challenges in Threat Information Sharing
    • Integrity checks: Further, many controllers also use integrity check mechanisms, typically based on CRCs or Checksums, to verify that programs have not changed  [30]. On the one hand, security products can verify that a specific program has been downloaded to the device by validating its signatures. On the other hand, CRC/Checksum integrity checks are not sufficient against threat actors who intentionally modify the PLC code  [32]. There is a need for a wider deployment of secured integrity signatures  [6].